1a) You purchased a share of stock for $35.40 seven years ago and just sold it today for $58.37. No

dividends were paid out over the seven years but you did receive an accumulated dividend of $5.80

when you sold the stock. What is your (a) Dollar gain or loss; (b) HPR; (c) Simple annual return; and (d)

Compound annual return?

1b) You bought a stock for $60.54 and sold it 5 months later for $68.95 with no dividends paid. What is

your (a) Dollar gain or loss; (b) HPR; (c) Simple annual return; and (d) Compound annual return?
Answer #1

1a.

Dollar Gain = (58.37 - 35.40) + 5.80 = $28.77

HPR = (58.77 - 35.40 + 5.80)/35.40 = 82.40%

Simple Return Rate = 0.8240/7 = 11.77%

CAGR = (1.8240)1/7 - 1 = 8.97%